# GITA: Graph to Visual and Textual Integration for Vision-Language Graph Reasoning Project Page: v-graph.github.io

### v4

@@ 3.3 Visual Graph Augmentation
Visual graphs generated for the same graph $G$ can be considered as a unique data augmentation technique. Building on the four graph-related image styles introduced in the graph visualizer part of Sec 3.2, we propose the following corresponding augmentation strategies: layout augmentation, node shape augmentation, node outline style augmentation, and edge thickness augmentation. Specifically, layout augmentation involves altering the layout styles while keeping all other settings constant. Similarly, by changing only the respective attributes, we can implement node shape augmentation, node outline style augmentation, and edge thickness augmentation. These four proposed augmentation strategies can facilitate studies on the importance of each in enhancing the graph reasoning abilities of VLM reasoners.
@@ 3.4 Training
Formally, given $I_{G}$, $Q_{G}^{T}$, and $A_{t}$, the VLM reasoner is trained by minimizing the negative log-likelihood as follows:
$\mathcal{L}_{\phi}=-\sum_{i=1}^{\lvert A\rvert}log\ p_{\phi}(A_{i}\mid F_{aligned},Q_{G}^{T},A_{<i}),$ (5)
where $\phi$ is trainable parameters and $A_{i}$ denotes the prediction token at the $i$-th position. Besides, $A_{<i}$ represents the first $i-1$ predicted tokens. During the inference process, GITA is capable of accepting structure graphs as inputs and performing graph reasoning in an end-to-end manner.
@@ 4.1 Construction
The GVLQA dataset is created by utilizing the graph visualizer, the graph describer, and the questioner in GITA to generate vision-language-based question-answer pairs for graph reasoning on an open-source graph dataset. Specifically, we first extract both the original graph structures and the ground-truth outputs from the NLGraph-full dataset (Wang et al., 2023). Then the graph visualizer (detailed in Sec 3.2) and the graph describer (outlined in Sec 3.2) are used to generate visual graphs and textual descriptions for these original graph structures, respectively. Afterward, the questioner (described in Sec 3.2) further improves and enriches the textual descriptions by converting them into textual queries. At the same time, it transforms the ground-truth output into text-based answers, following specific output requirements. By combining these visual graphs, textual queries, and text-based answers, we obtain the Graph-based Vision-Language Question Answering (GVLQA) dataset.
In the process of establishing GVLQA, we employed graphviz (Gansner & North, 2000) to instantiate the graph visualizer. This choice is made due to its multitude of pre-defined layout algorithms, which enable convenient adjustment of visual graph layouts. Additionally, manual-template-based constructed queries are utilized as the questioner because these tasks are famous with well-defined requirements.
@@ 4.2 Structure
The GVLQA dataset comprises 526K samples, each consisting of a visual graph, a textual query, and the corresponding answer. It is divided into five subsets: GVLQA-BASE, and four augmentation subsets GVLQA-AUGLY, GVLQA-AUGNS, GVLQA-AUGNO, and GVLQA-AUGET. In GVLQA-BASE, the visual graphs are uniformly styled. The remaining four augmentation subsets are derived from GVLQA-BASE through the four visual graph augmentations (Sec 3.3), varying in six different layouts, three node shapes, four node outline styles, and four degrees of edge thickness, respectively. Detailed statistics for each subset are presented in Table 4 in Appendix B.
@@
  • Connectivity (Sedgewick, 2001) (denoted by Connect): to determine whether two randomly selected nodes $u$ and $v$ in an undirected graph are connected.
  • Cycle (Sedgewick, 2001): to identify whether a cycle exists in an undirected graph.
  • Topological Sort (Kahn, 1962) (denoted by TS): to find a valid topological sorting for a directed acyclic graph. Here a topological sorting is to output a linear ordering of nodes such that for every directed edge $u\leftarrow v$, node $u$ comes before $v$ in the ordering.
  • Shortest Path (Dijkstra, 1959) (denoted by SP): to find the shortest path between two nodes in a weighted undirected graph, where the shortest path between two nodes is the path connecting the two nodes with the minimum sum of edge weights along the path.
  • Maximum Flow (Ford & Fulkerson, 1956) (denoted by MaxFlow): to calculate the maximum flow from a source node to a sink node in a network graph.
  • Bipartite Graph Matching (Karp et al., 1990) (denoted by BGM): to find a matching set in a bipartite graph that has the largest number of edges, where a matching set is a collection of edges in which no two edges share any common node.
  • Hamilton Path (Gould, 2003) (denoted by HP): to find a valid Hamilton path in an undirected graph, where a Hamiltonian path is a path that traverses each node in a graph exactly once.
Figure 6 offers illustrations for those tasks in the GVLQA-BASE dataset. Illustrations of all GVLQA subsets are provided in Appendix G.
@@ 5 Experiments
In this section, we extensively evaluate the performance of LLM baselines and the proposed GITA on the GVLQA-BASE and five real-world datasets. To better clarify the reasoning capabilities of solely visual graphs, we also test GITA without the textual descriptions of graphs, which can be considered as a variant of GITA and denoted as vision-only (VO). In this case, the visual graph is the only information source for graph reasoning. Additionally, we investigate the importance of visual graph augmentation (Sec 3.3) strategies, by comparison among GITA-7B trained on GVLQA-BASE and on the other augmentation subsets of GVLQA (Sec 4.2). Lastly, we investigate the effectiveness of using GVLQA as the pretrained dataset on real-world datasets. The evaluation metrics for all experiments are accuracy by exact matching. For the fine-tuning setting, we fine-tune the LoRA adapters (Hu et al., 2021) for all weight matrices within the text decoder of the VLM reasoner, while keeping the vision encoder in the VLM reasoner frozen. More detailed experimental settings are in Appendix F.
@@ 5.1 Evaluation on the GVLQA-BASE Dataset
In this subsection, we perform experiments on the GVLQA-BASE dataset by comparing GITA with popular LLMs-based baselines including GPT-4 Turbo (OpenAI, 2023), LLaMA2-7B/13B (Touvron et al., 2023), and Vicuna-7B/13B (Zheng et al., 2023), under both zero-shot and fine-tuning settings. The experimental results are shown in Table 1. From these results, we summarize the following observations.
@@
Observation 1: GITA outperforms LLM Baselines. As can be seen in Table 1, GITA consistently outperforms the LLM baselines under the same setting. This underscores its SOTA effectiveness in instruction-based graph reasoning tasks, showing robust capabilities across different parameter scales under both fine-tuning and zero-shot settings. Moreover, under the fine-tuning setting, incorporating the vision modality consistently benefits 7B models. But for 13B models, the performance of some tasks may degrade. This could be attributed to the greater challenge of aligning the representation of visual and textual modalities in larger 13B models compared to 7B models, in the case of only fine-tuning LoRA adapters within the text decoder. We speculate that full training could potentially address this issue. However, we leave this as future work due to resource constraints.
@@
Observation 2: Mainstream open-source VLM/LLMs lack fundamental graph reasoning abilities. The zero-shot results illustrate that prominent open-source open-source LLMs or VLMs, including LLaMA2, Vicuna and LLaVA, exhibit minimal graph reasoning capabilities on the GVLQA-BASE dataset. Specifically, these models produce random answers, randomly responding with either "Yes." or "No." for tasks involving Connect and Cycle, resulting in a performance close to 50%. In contrast, current SOTA closed-source LLMs or VLMs, including GPT-4 Turbo and GPT-4V, demonstrate superior zero-shot performance compared to the aforementioned open-source models. This observation implies that current open-source LLMs and VLMs lack of basic graph reasoning ability, which may be attributed to the insufficient availability of relevant training data. Such observation also enhances the motivation for us to propose the GVLQA dataset, with the aim of improving the graph reasoning capabilities of VLMs/LLMs.
@@
Observation 3: Model size increase enhances fine-tuning graph reasoning capabilities. The comparison of VLMs/LLMs with different parameter sizes, specifically 7B-parameter models and 13B-parameter models, highlights the benefits of increasing the model size for graph reasoning capabilities. In this regard, GITA-13B outperforms its counterpart with 7B parameters (GITA-7B) both on average and across four out of seven tasks. However, it is worth noting that GITA-13B does not outperform GITA-7B in the other three tasks. We hypothesize that this discrepancy may be attributed to insufficient modality alignment due to LoRA-only fine-tuning.
@@
Observation 4: Vision and Text Modalities proficient in different types of Graph Reasoning Tasks. We explore the individual capabilities of the visual and textual modalities within the GITA framework. The results indicate that the text and vision modalities can complement each other and contribute to better performance than individual ones, as removing either modality leads to a drop in performance in most cases (Vicuna & GITA (VO) and GPT-4 Turbo & GITA (VO) in Table 1). While the graph reasoning capability provided by the vision may not be as strong as that of the text modality in most cases, relying solely on vision still enables the model to possess basic graph reasoning abilities. Specifically, the model outperforms text-based LLMs in 2/7 tasks (Cycle and BGM) when relying solely on vision. This consistent improvement across all comparison groups demonstrates the potential of the vision modality to excel in certain graph reasoning tasks, leveraging its ability to capture visual patterns like cycles and graph properties such as bipartition. In contrast, text exhibits a higher proficiency than vision modality in sequence-related graph reasoning problems, particularly in tasks like TS, SP, and HP, which require constructing ordered node sequences.
@@ 5.2 Evaluation for the Visual Graph Augmentations
To assess the impact of proposed visual graph augmentation strategies, including layout, node shape, node outline style, and edge thickness augmentations, we compare the performance of vision-only GITA-7B models trained on four augmented subsets of GVLQA and on GVLQA-BASE (without augmentation). The comparative results are presented in Table 2. To fully utilize the visual information in visual graphs, we fine-tune the visual encoder of VLMs in addition to the vision-to-text projector and the LoRA adapters within the text decoder in this part.
@@
From the obtained results, a significant enhancement in overall performance is observed with the introduction of layout augmentation (GVLQA-AUGLY). The average performance improves remarkably from 38.91% to 63.36%. Notably, significant improvements were observed in SP (5.72% to 76.55%), HP (1.11% to 70.74%), and Connect (59.97% to 87.18%). These findings highlight the critical role of layout augmentation in the generation of visual graphs. In other words, this observation suggests the potential for creating larger-scale datasets for vision-language-based graph reasoning, which could significantly contribute to the advancement of this field. Conversely, the other three augmentations do not yield such substantial performance improvements, further emphasizing the importance of layout augmentation in vision-language-based reasoning.
5.3 Evaluation on the Real-World Datasets
@@
In this section, we illustrate the effectiveness of GITA on the ca-GrQC Leskovec et al. (2007) and ca-HepTh Leskovec et al. (2007) datasets for the link prediction task, and on the PolBlog Adamic & Glance (2005), Cora Yang et al. (2016) and CiteSeer Yang et al. (2016) datasets for the node classification task. Table 5 in the appendix presents the statistics of those datasets, where graph data could involve thousands of nodes/edges, making it infeasible to feed the entire graph into the model. Consequently, we employ $k$-hop subgraph sampling discussed in Sec 3.2, specifically with $k=2$, to satisfy the token length restriction of LLMs and visual graph scope effectively.
@@
The experimental Results are shown in Table 3. The presented results have also verified the effectiveness of GITA and the proposed layout augmentation. Notably, we highlight the benefits of using GVLQA-BASE as pretrained dataset by comparing it with GITA-7b. Performance improvements of 0.67% and 0.92% were observed under zero-shot and fine-tuning settings, respectively. This underscores the potential application value of the proposed GVLQA dataset.

### v5

@@ 3.3 Visual Graph Augmentation
Visual graphs generated for the same graph $G$ can be considered as an unique data augmentation technique. Building on the four graph-related image styles introduced in the graph visualizer part of Sec 3.2, we propose the following augmentation strategies: layout augmentation, node shape augmentation, node outline style augmentation, and edge thickness augmentation. Specifically, layout augmentation involves altering the layout styles while keeping all the other settings constant. Similarly, by changing only the respective attributes, we can implement node shape augmentation, node outline style augmentation, and edge thickness augmentation. These four proposed augmentation strategies facilitate studies on the importance of each in enhancing the graph reasoning abilities of VLM reasoners.
@@ 3.4 Training
Formally, given $I_{G}$, $Q_{G}^{T}$, and $A_{t}$, the VLM reasoner is trained by minimizing the following negative log-likelihood:
$\mathcal{L}_{\phi}=-\sum_{i=1}^{\lvert A\rvert}\log\ p_{\phi}(A_{i}\mid F_{aligned},Q_{G}^{T},A_{<i}),$ (5)
where $\phi$ is the trainable parameter and $A_{i}$ denotes the prediction token at the $i$-th position. Besides, $A_{<i}$ represents the first $i-1$ predicted tokens. During the inference process, GITA is capable of accepting structure graphs as inputs and performing graph reasoning in an end-to-end manner.
@@ 4.1 Construction
The GVLQA dataset is created by utilizing the graph visualizerthe graph describer, and questioner in GITA to generate vision-language-based question-answer pairs for graph reasoning on an open-source graph dataset. Specifically, we first extract both the original graph structures and the ground-truth outputs from the NLGraph-full dataset [59]. Then the graph visualizer (detailed in Sec 3.2) and the graph describer (outlined in Sec 3.2) are used to generate visual graphs and textual descriptions for these original graph structures, respectively. Afterwards, the questioner (described in Sec 3.2) further improves and enriches the textual descriptions by converting them into textual queries. At the same time, it transforms the ground-truth output into text-based answers, following specific output requirements. By combining these visual graphs, textual queries, and text-based answers, we obtain the Graph-based Vision-Language Question Answering (GVLQA) dataset.
In the process of establishing GVLQA, we employed graphviz [17] to instantiate the graph visualizer. This choice is made due to its multitude of pre-defined layout algorithms, which enable convenient adjustment of visual graph layouts. Additionally, manual template-based constructed queries are utilized as the questioner because these tasks are famous with well-defined requirements.
@@ 4.2 Structure
The GVLQA dataset comprises 526K samples, each consisting of a visual graph, a textual query, and its corresponding answer. It is divided into five subsets: GVLQA-BASE, and four augmentation subsets GVLQA-AUGLY, GVLQA-AUGNS, GVLQA-AUGNO, and GVLQA-AUGET. In GVLQA-BASE, the visual graphs are uniformly styled. The remaining four augmentation subsets are derived from GVLQA-BASE through the four visual graph augmentations (Sec 3.3), varying in six different layouts, three node shapes, four node outline styles, and four degrees of edge thickness, respectively. Detailed statistics of the four subsets are shown in Table 6 of Appendix C.
@@
  • Connectivity [52] (denoted Connect): Determine whether two randomly selected nodes $u$ and $v$ in an undirected graph are connected.
  • Cycle [52]: Identify whether a cycle exists in an undirected graph.
  • Topological Sort [35] (denoted TS): Find a valid topological sort for a directed acyclic graph. Here, topological sort outputs a linear ordering of the nodes such that for every directed edge $u\leftarrow v$, node $u$ comes before $v$ in the ordering.
  • Shortest Path [13] (denoted SP): Find the shortest path between two nodes in a weighted undirected graph. The shortest path between two nodes is the path connecting the two nodes with the minimum sum of edge weights along the path.
  • Maximum Flow [15] (denoted MaxFlow): Calculate the maximum flow from a source node to a sink node in a network graph.
  • Bipartite Graph Matching [36] (denoted BGM): Find a matching set in a bipartite graph with the largest number of edges. A matching set is a collection of edges in which no two edges share any common node.
  • Hamilton Path [20] (denoted HP): Find a valid Hamilton path in an undirected graph. A Hamiltonian path is a path that traverses each node in a graph exactly once.
Figure 6 offers illustrations for these tasks in the GVLQA-BASE dataset. Illustrations of all the GVLQA subsets are provided in Appendix H.
@@ 5 Experiments
In this section, we extensively evaluate the performance of LLM baselines and the proposed GITA on the GVLQA-BASE and five real-world datasets. To better clarify the reasoning capabilities of solely visual graphs, we also test GITA without the textual descriptions of graphs, which can be considered as a variant of GITA and denoted as vision-only (VO). In this case, the visual graph is the only information source for graph reasoning. Additionally, we investigate the importance of visual graph augmentation (Sec 3.3) strategies, by comparison GITA-7B trained on GVLQA-BASE and on the other augmentation subsets of GVLQA (Sec 4.2). Lastly, we investigate the effectiveness of using GVLQA as the pretrained dataset on real-world datasets. The evaluation metrics for all experiments are accuracy by exact matching. For the fine-tuning setting, we fine-tune the LoRA adapters [28] for all weight matrices in the text decoder of the VLM reasoner, while keeping the vision encoder in the VLM reasoner frozen. More detailed experimental settings are in Appendix G.
@@ 5.1 Evaluation on the GVLQA-BASE Dataset
In this subsection, we perform experiments on the GVLQA-BASE dataset to compare GITA with popular LLMs including GPT-4 Turbo [49], LLaMA2-7B/13B [57], and Vicuna-7B/13B [73], under both zero-shot and fine-tuning settings. The experimental results are shown in Table 1. Based on these results, we can obtain the following observations.
@@
Observation 1: GITA Outperforms LLM Baselines. As can be seen in Table 1, GITA consistently outperforms the LLM baselines under the same setting. This underscores its SOTA effectiveness in instruction-based graph reasoning tasks, showing robust capabilities across different parameter scales under both fine-tuning and zero-shot settings. Moreover, under the fine-tuning setting, incorporating the vision modality consistently benefits 7B models. But for the 13B models, the performance of some tasks may degrade. This could be attributed to the greater challenge of aligning representations of the visual and textual modalities in the larger 13B models compared to the 7B models, in the case of only fine-tuning LoRA adapters in the text decoder. We speculate that full training could potentially address this issue. However, we leave this as future work due to resource constraints.
@@
Observation 2: Mainstream Open-source VLM/LLMs Lack Fundamental Graph Reasoning Abilities. The zero-shot results illustrate that prominent open-source LLMs or VLMs, including LLaMA2, Vicuna, and LLaVA, exhibit minimal graph reasoning capabilities on the GVLQA-BASE dataset. Specifically, these models produce random answers, i.e., randomly responding with either "Yes." or "No." for tasks involving Connect and Cycle, resulting in a performance close to 50%. Current SOTA closed-source LLMs or VLMs, including GPT-4 Turbo and GPT-4V, demonstrate superior zero-shot performance compared with the aforementioned open-source models. This observation implies that current open-source LLMs and VLMs lack basic graph reasoning ability, which may be attributed to the insufficient availability of relevant training data. Such observation also enhances our motivation to propose the GVLQA dataset, with the aim of improving the graph reasoning capabilities of VLMs/LLMs.
@@
Observation 3: Increasing Model Size Leads to Better Graph Reasoning Capabilities. The comparison of VLMs/LLMs with different parameter sizes, specifically 7B and 13B models, verify the benefits of increasing the model size for graph reasoning capabilities. In this regard, GITA-13B outperforms its counterpart with 7B parameters (GITA-7B) both on average and across four of the seven tasks. However, it is worth noting that GITA-13B does not outperform GITA-7B on the other three tasks. We hypothesize that this discrepancy may be attributed to insufficient modality alignment due to LoRA fine-tuning.
@@
Observation 4: Vision and Text Modalities Proficient in Different Types of Graph Reasoning Tasks. We explore the individual capabilities of the visual and textual modalities within the GITA framework. The results indicate that the text and vision modalities can complement each other and contribute to better performance than individual ones, as removing either modality leads to performance drops in most cases (Vicuna & GITA (VO) and GPT-4 Turbo & GITA (VO) in Table 1). While the graph reasoning capability provided by the vision modality may not be as strong as that of the text modality in most cases, relying solely on vision still enables the model to possess basic graph reasoning abilities. Specifically, the model outperforms text-based LLMs in 2 of the 7 tasks (Cycle and BGM) when relying solely on vision. This consistent improvement across all comparison groups demonstrates the potential of the vision modality to excel in certain graph reasoning tasks, leveraging its ability to capture visual patterns like cycles and graph properties such as bipartition. In contrast, text exhibits a higher proficiency than vision modality in sequence-related graph reasoning problems, particularly on tasks such as TS, SP, and HP, which require constructing ordered node sequences.
@@ 5.2 Evaluation for the Visual Graph Augmentations
To assess the impact of the proposed visual graph augmentation strategies (including layout, node shape, node outline style, and edge thickness augmentations), we compare the performance of vision-only GITA-7B models trained on the four augmented subsets of GVLQA and on GVLQA-BASE (without augmentation). The results are presented in Table 2. To fully utilize the visual information in visual graphs, we fine-tune the visual encoder of VLMs in addition to the vision-to-text projector and the LoRA adapters within the text decoder in this experiment.
@@
As can be seen from the results, a significant enhancement in overall performance is observed with the introduction of layout augmentation (GVLQA-AUGLY). The average performance improves remarkably from 38.91% to 63.36%. Notably, significant improvements are observed on SP (5.72% to 76.55%), HP (1.11% to 70.74%), and Connect (59.97% to 87.18%). These findings highlight the critical role of layout augmentation in generating visual graphs. In other words, this observation suggests the potential for creating larger-scale datasets for vision-language-based graph reasoning, which could significantly contribute to advancing this field. Conversely, the other three augmentations do not yield such substantial performance improvements, further emphasizing the importance of layout augmentation in vision-language-based reasoning.
5.3 Evaluation on Real-World Datasets
@@
In this section, we study the effectiveness of GITA on the ca-GrQC [42] and ca-HepTh [42] datasets for the link prediction task, and on the PolBlog [1], Cora [65] and CiteSeer [65] datasets for the node classification task. Table 8 in the appendix C presents the statistics of these datasets. The graph can have thousands of nodes/edges, making it infeasible to feed the entire graph into the model. Consequently, we employ $k$-hop subgraph sampling (with $k=2$) discussed in Sec 3.2 to satisfy the token length restriction of LLMs and visual graph scope effectively.
@@
The experimental results are presented in Table 3. It is evident that GITA consistently outperforms the LLM baselines, and its performance progressively improves with the addition of layout augmentation and the use of the GVLQA checkpoint. Notably, we emphasize the advantages of using GVLQA-BASE as a pretrained dataset by comparing it with GITA-7b. Performance improvements of 0.67% and 0.92% are observed in the zero-shot and fine-tuning settings, respectively. This highlights the potential application value of the proposed GVLQA dataset.
5.4 Comparison of GITA with Dedicated Graph Baselines
Table 4: Accuracy (%) comparisons among dedicated GNNs and GITAs on GVLQA-Base.
Connect Cycle TS SP MaxFlow BGM HP Avg
GCN 79.65 70.89 45.71 44.56 56.44 76.70 32.22 58.02
SAGE 82.72 73.58 44.51 49.25 50.67 81.00 36.67 59.78
GITA-7B 98.95 96.67 41.12 32.15 20.00 93.19 29.26 58.76
GITA-13B 99.14 95.60 38.69 40.47 20.66 92.12 33.33 60.00
Though GITA is designed for language-based general graph reasoning settings, which are much more user-friendly (by user-readable natural language) and general (unique model architecture for various scenarios) than the typical application of dedicated GNNs, it remains essential to conduct a comprehensive comparison with specialized GNNs to elucidate the strengths and limitations of GITA’s applicability and capabilities. To this end, we assess the graph reasoning abilities of GITA against dedicated GNNs, including GCN [38] and SAGE [24], using the GVLQA-Base dataset, as detailed in Table 4. In addition, we explore and compare the effects of $k$-hop subgraph sampling on the proposed GITA and GNN baselines. Using the ca-Hepth dataset, we analyze the impact of increasing the number of hops $k$ on the reasoning time and performance of both GITA and GNNs. The results are in Table 5.
Overall Graph Reasoning Ability Comparison. As shown in Table 4, compared to the dedicated GNNs, the fine-tuned GITA-7B models have comparable average graph reasoning performance, with the larger GATA-13B model performs slightly better. In particular, compared to GNNs, the GITA model shows a stronger ability in recognizing local structures in the graphs (Connect and Cycle) and to accomplish tasks with obvious layout heuristics (BGM). We believe that this advantage comes from GITA’s visual perception. For SP and MaxFlow, GITA’s performance is inferior to GNNs. This may be because GNNs process edge weight more effectively through the message-passing mechanism.
Scalability and Performance Variation with Different Numbers of Hops $k$. The inference time results are shown in Table 5. As can be seen, GITA demonstrates inferior scalability compared to the GNN baselines. Its scalability remains stable as the sampled graph size (i.e., $k$) increases.
Table 5: Accuracy (%) and Inference Time (in parentheses) for GNNs and GITA on ca-Hepth Dataset with different subgraph sampling hop number $k\in\{1,2,3,4\}$.
GCN SAGE GITA-7B
k=1 93.27 (0.02s) 94.40 (0.03s) 90.33 (17.23min)
k=2 94.49 (0.04s) 94.43 (0.04s) 91.13 (17.66min)
k=3 91.10 (0.04s) 90.95 (0.18s) 90.31 (17.22min)
k=4 81.92 (0.05s) 83.60 (0.22s) 86.10 (17.01min)
From the accuracy results in Table 5, GITA, GCN, and SAGE achieve their peak performance at $k=2$, suggesting that a small sampled graph size suffices for optimal performance. Though the dedicated GNNs attain higher peak performance than GITA, they exhibit performance declines as $k$ increases (e.g., 3 or 4), while GITA’s performance is more stable w.r.t. $k$.